OpenStudy (anonymous):

Help @jim_thompson5910

OpenStudy (anonymous):

\[\sec \theta \cos \theta \] Simplify

jimthompson5910 (jim_thompson5910):

Hint: sec(theta) = 1/cos(theta)

OpenStudy (anonymous):

so would it be 1?

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

Use the unit circle to find the inverse function value in degrees.\[\tan^{-1} \sqrt{3}\]

jimthompson5910 (jim_thompson5910):

use this unit circle http://3.bp.blogspot.com/-81sNFZa1qWw/UMhdT_4s7nI/AAAAAAAAAgA/oGy2QtxD8Q8/s1600/Unit+circle-tan1.JPG the values to the right of the parenthesis are the tangent values

jimthompson5910 (jim_thompson5910):

so, for example, tan(pi/4) = 1

jimthompson5910 (jim_thompson5910):

since pi/4 radians = 45 degrees, we can say tan(45) = 1

OpenStudy (anonymous):

so would it be 120 degrees? i dont really understand this very well

jimthompson5910 (jim_thompson5910):

look in the first quadrant

jimthompson5910 (jim_thompson5910):

which is in the upper right

jimthompson5910 (jim_thompson5910):

where do you see \(\large \sqrt{3}\) after the parenthesis ?

OpenStudy (anonymous):

30 degrees?

jimthompson5910 (jim_thompson5910):

open that image up in paint and circle it for me

jimthompson5910 (jim_thompson5910):

perfect, thank you

jimthompson5910 (jim_thompson5910):

what is the angle that corresponds to this point

OpenStudy (anonymous):

30

jimthompson5910 (jim_thompson5910):

the angle is off to the left of the parenthesis

OpenStudy (anonymous):

im not sure if i understand

jimthompson5910 (jim_thompson5910):

see in the parenthesis where it says 1/2, sqrt(3)/2

OpenStudy (anonymous):

okay yeah

jimthompson5910 (jim_thompson5910):

off to the left is what

OpenStudy (anonymous):

\[\pi / 3\]

jimthompson5910 (jim_thompson5910):

that's your angle in radians

jimthompson5910 (jim_thompson5910):

convert that to degrees to get your answer

jimthompson5910 (jim_thompson5910):

to do so, you multiply by 180/pi

OpenStudy (anonymous):

60

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

for the next two we are trying to figure out what values for \[\theta(0\le \theta \le2\pi) \] satisfy the equation.

OpenStudy (anonymous):

1) \[3\sin \theta = \sin \theta - 1\] 2) \[\tan^2 \theta = -3/2 \sec \theta \]

jimthompson5910 (jim_thompson5910):

\[\large 3\sin \theta = \sin \theta - 1\] \[\large 3\sin \theta - \sin \theta = - 1\] \[\large 2\sin \theta = - 1\] \[\large \sin\theta = -\frac{1}{2}\] I'll let you finish

OpenStudy (anonymous):

i dont think i got it.. i ggot 57.296

jimthompson5910 (jim_thompson5910):

use the unit circle, when is sine equal to -1/2 ?

OpenStudy (anonymous):

\[-\sqrt{3}\] ?

jimthompson5910 (jim_thompson5910):

look for points that have a y coordinate of -1/2

jimthompson5910 (jim_thompson5910):

then tell me the angles that correspond to these points

OpenStudy (anonymous):

okay i figured out the answer to that one.. let me see if i can try and figure out the second one... \[\tan^2\theta - \sec \theta = -3/2\] is that right?